find the general integral of the equation (x-y)p+(y-x-z)q=z and a particular solution through the circle z=1,x^2+y^2=1
1
Expert's answer
2020-05-19T08:43:02-0400

By Lagrange’s method the auxiliary equations of the given differential equation are as following:

"\frac{𝑑𝑥}{𝑥−𝑦}=\frac{𝑑𝑦}{𝑦−𝑥−𝑧}=\frac{𝑑𝑧}{𝑧}"

"\implies \frac{𝑑𝑥}{𝑥−𝑦}=\frac{𝑑𝑦}{𝑦−𝑥−𝑧}=\frac{𝑑𝑧}{𝑧} = \frac{dx+dy+dz}{0} \\
\implies dx+dy+dz =0 \implies x+y+z = c_1"

Hence from the first fraction, we have "\frac{𝑑𝑥}{𝑥−𝑦}=\frac{𝑑𝑦}{𝑦−𝑥−𝑧}"

"\implies \frac{𝑑𝑦}{𝑦−c_1+y}=\frac{𝑑z}{z}"

Now, by integrating on both sides we get "(1/2)\log(2y-c_1) =\log(z)+\log(\sqrt{c_2})
\implies c_2 = \frac{y-x-z}{z^2}" .


Since "x+y+z=c_1" and "\frac{y-x-z}{z^2}=c_2" are independent solutions of the given differential equation, so a solution of the given differential equation is

"c_2= f(c_1) \implies \frac{y-x-z}{z^2} = f(x+y+z)" ________(1)


On "x^2+y^2 = 1, z=1" , so put "x= \cos{(t)}, y = \sin{(t)}" and put in solution, we get

"\sin{(t)}-\cos{(t)}=f(\cos{(t)}+\sin{(t)})" ________(2)

So, the solution of the given differential equation is given by (1), where the function is given by (2).